# Domain Generalization and Adaptation in Intensive Care with Anchor Regression

### v1

@@ Abstract
The performance of predictive models in clinical settings often degrades when deployed in new hospitals due to distribution shifts. This paper presents a large-scale study of causality-inspired domain generalization on heterogeneous multi-center intensive care unit (ICU) data. We apply anchor regression and introduce anchor boosting, a novel, tree-based nonlinear extension, to a large dataset comprising 400,000 patients from nine distinct ICU databases. The anchor regularization consistently improves out-of-distribution performance, particularly for the most dissimilar target domains. The methods appear robust to violations of theoretical assumptions, such as anchor exogeneity. Furthermore, we propose a novel conceptual framework to quantify the utility of large external data datasets. By evaluating performance as a function of available target-domain data, we identify three regimes: (i) a domain generalization regime, where only the external model should be used, (ii) a domain adaptation regime, where refitting the external model is optimal, and (iii) a data-rich regime, where external data provides no additional value.
@@ 1 Introduction
A standard assumption in predictive modeling is that training and test data come from the same distribution. This assumption often fails in real-world scenarios. For example, in clinical applications, test data may originate from a different time period or hospital than the training data. When these distribution shifts occur, model performance tends to drops significantly (Barak-Corren et al., 2021; Guo et al., 2021; Roland et al., 2022; Yang et al., 2022; van de Water et al., 2024; Hüser et al., 2024).
@@
In this paper, we apply and extend anchor regression for medical prediction. Recognizing that linear models may be insufficient to capture the complex feature interactions in clinical data, we first propose a novel, nonlinear extension to anchor regression based on gradient boosting trees (Friedman, 2001), popular in clinical predictive modeling (Hyland et al., 2020; Lyu et al., 2024; Hüser et al., 2024). We then conduct a large-scale empirical study, applying linear anchor regression and our nonlinear extension to predict adverse events in intensive care units (ICUs). We assess the method’s ability to improve distributional robustness on a strongly heterogeneous dataset aggregating nine ICU databases (Burger et al., 2024, 2025). Finally, we propose a conceptual framework for quantifying the utility of large external datasets, particularly when target domain data is scarce.
@@ 1.1 Our contribution
We conduct the largest-scale application of anchor regression (Rothenhäusler et al., 2021) and its variants to date, using a dataset of 400,000 patients and 10 million observations from nine distinct ICU databases. To our knowledge, this work is also the first application of anchor regression and the largest application of a causality-inspired method to a medical prediction problem. In a setting where many other domain generalization methods have been shown to provide little benefit over simple baselines (Rockenschaub et al., 2024), we show that the anchor methods yield significant performance improvements, particularly for the most out-of-distribution target domains.
@@ 2 Data description
We use data from Burger et al. (2024, 2025) who build upon the R-package ricu (Bennett et al., 2023) to harmonize and aggregate ICU data from different sources. We will describe this dataset in further detail next.
@@ 2.1 ICU datasets included
We focus on the following 9 ICU datasets: The eICU Collaborative Research Database (Pollard et al., 2018) is a multi-center critical care database of 207207207 hospitals throughout the USA. It contains data from 188’257 patient stays from 2015 and 2016. The Medical Information Mart for Intensive Care (MIMIC) III (Johnson et al., 2016) and IV (Johnson et al., 2023) contain data from the Beth Israel Deaconess Medical Center (BIDMC) in Boston (USA). The BIDMC switched critical care information systems in 2008 from Philips CareVue to iMDsoft Metavision. We consider only the CareVue subset of MIMIC-III to avoid an overlap with MIMIC-IV. This MIMIC-III CareVue subset contains data from 34’154 patient stays from 2001–2008 and includes neonatal patient stays. MIMIC-IV contains data from 93’679 patient stays from 2008–2022. The Northwestern ICU database (NWICU) (Moukheiber et al., 2024) is a multi-center critical care database of 12 hospitals around Chicago (USA). It contains data from 28’150 patient stays from 2020–2022. The High-Resolution ICU dataset (HiRID) (Hyland et al., 2020), Amsterdam University Medical Center database (AUMCdb) (Thoral et al., 2021), and Salzburg Intensive Care database (SICdb) (Rodemund et al., 2024), are single-center European datasets. HiRID contains data from 33’586 patient stays from 2008–2016. AUMCdb contains data from 22’897 patient stays from 2003–2016. SICdb contains data from 27’115 patient stays from 2013–2021. The Paediatric Intensive Care database (PICdb) (Li et al., 2019) contains data from the Children’s Hospital of Zhejiang University School of Medicine in China. It contains data from 13’516 patient stays from 2010–2018. The Critical Care Database Comprising Patients With Infection at Zigong Fourth People’s Hospital (Xu et al., 2022) in China contains data of patients with a suspected infection. It contains data from 2’583 patient stays from 2019 and 2020. All datasets except for AUMCdb are available on PhysioNet (Goldberger et al., 2000). We summarize these data in table˜2 in appendix˜A.
@@ 2.3 Patient outcomes
We visualize the resulting outcome distributions in figure˜1 and provide summary statistics in table˜2 in appendix˜A. See also github.com/eth-mds/icu-features.
@@ 2.4 Feature engineering and pre-selection of variables
The hourly time-series data, extracted by Burger et al. (2024, 2025) using ricu (Bennett et al., 2023) must be transformed into a static feature set to be used with standard linear or tree-based methods. Following the literature, we engineer features over a backwards looking time window to capture some temporal dynamics of a patient’s history. We use an 8 hour horizon for circulatory failure and log(lactate) prediction and a 24 hour horizon for acute kidney injury and log(creatinine) prediction. These features include a missingness indicator, the mean, standard deviation, maximum, minimum, and a linear trend for continuous variables, the mode and a missingness indicator for categorical variables, and an indicator and the average rate for treatments. See github.com/eth-mds/icu-features and section˜A.2 for details.
For computational reasons and to reduce model complexity, we do not use all available variables of Burger et al. (2024, 2025)’s export in our models. For the prediction of circulatory failure and log(lactate), we use the top 20 variables of Hyland et al. (2020) according to their table 1. For the prediction of akute kidney injury and log(creatinine), we use the top variables of Lyu et al. (2024) according to their figure 8a. See section˜A.3 for details. Together with feature engineering, this results in 100–200 covariates for each task.
@@ 2.5 Sources of heterogeneity
Because of the cohort selection and variable availability, expect the largest differences in cohort composition and variable availability for NWICU, PICdb, and Zigong. We therefore designate these as “truly out-of-distribution” datasets. In section˜4, except for sections˜4.6 and 4.7, we will use these only for evaluation and not for training.
@@ 3 Methods
This section details our methodological approach. We begin by introducing linear anchor regression (Rothenhäusler et al., 2021). We then propose a anchor boosting, a novel, nonlinear extension based on gradient boosting trees. Finally, we describe our refitting procedure and introduce a taxonomy to quantify the value of external data.
@@ 3.1 Linear anchor regression
Consider a discrete anchor variable given by environments $e\in{\cal E}$, for example, the ICU datasets =${\cal E}=\{\mathrm{AUMCdb},\ \mathrm{eICU},\ \ldots,\ \mathrm{SICdb}\}$. The linear anchor regression estimator is defined as
@@
where $\gamma\geqslant 1$ is the invariance regularization parameter and $n_{e}$ are the number of samples from environment $e$. That is, anchor regression penalizes differences in the environments’ mean residuals. From a theoretical perspective, the anchor regression estimator optimizes a worst-case risk over a set new, unseen environments $e\notin{\cal E}$, for example $e=\mathrm{PICdb}$, with distribution shifts similar to the heterogeneity seen within the training data, but of a larger magnitude scaled by $\gamma$.
@@ 3.2 Nonlinear anchor boosting (regression)
The theoretical robustness guarantees of anchor regression mentioned above apply for a linear model (Rothenhäusler et al., 2021) or when using a nonlinear embedding (Šola et al., 2025). This does not prevent us from optimizing equations˜1 and 2, replacing $X_{i}\beta$ with a nonlinear function $f(X_{i})$. We focus here on boosted tree learners (Friedman, 2001), a popular method for ICU data (Hyland et al., 2020; Lyu et al., 2024). Analogously to (2), let
@@
be the anchor loss with gradient and Hessian
[…]
@@
A nonlinear extension of anchor regression based on boosting was already proposed by Bühlmann (2020). However, this proposal does not use second order optimization to update the leaf node values, which appears to be crucial when using larger values of $\gamma$, especially for classification as described in section˜3.3. See section˜B.3 for details. In related work, Ulmer et al. (2025) fit a random forest to a linear rescaling of the data. In Ulmer and Scheidegger (2025) they discuss how this can be applied to anchor regression, resulting in split points for each tree chosen to directly minimize the anchor loss. However, this approach increases computational complexity beyond a near-linear relationship with sample size, making it prohibitive for the large datasets that we address in this paper.
@@ 3.3 Nonlinear anchor boosting (classification)
We apply anchor regularization to binary classification tasks. Kook et al. (2022) suggest to use the gradient of the log-likelihood as score residuals. We use a probit link function, as, in contrast to the logistic link, the resulting anchor classification objective is convex. See section˜B.2 for details.
Write $\Phi$ and $\varphi$ for the Gaussian distribution’s cumulative distribution function and probability density function. For binary classification with scores $f\in{\mathbb{R}}^{n}$ and $y\in\{-1,1\}^{n}$, the negative log-likelihood is $-\sum_{i=1}^{n}\log(\Phi(y_{i}f_{i}))$ with negative gradient $r:=y\cdot\varphi(f)/\Phi(yf)$. We thus apply the same procedure as in section˜3.2 to the loss
@@
replacing the squared error $\frac{1}{2}(y_{i}-f_{i})^{2}$ in equation˜3 with the negative log-likelihood $-\log(\Phi(y_{i}f_{i}))$ and the residuals $y_{i}-f_{i}$ with the likelihood score residuals $r_{i}$. We initialize with $\hat{f}^{0}=\Phi^{-1}(\frac{1}{n}\sum_{i=1}^{n}\frac{y_{i}+1}{2})$. See section˜B.1 for derivations of the gradient and Hessian of equation˜4.
@@ 4.1 Anchor regularization improves generalization to some ICU datasets
Figure 3: Anchor boosting’s OOD MSE predicting log(creatinine) as a function of $\gamma$, using one-hot-encoded dataset ID as anchor.
@@
Figure˜17 in appendix˜C shows anchor boosting’s OOD area under the precisision-recall curve (AuPRC) predicting acute kidney injury within 48 hours, the classification task corresponding to log(creatinine) regression. AuPRC is a common metric for predicting rare events on ICU data. Anchor regularization with $\gamma>1$ yields a considerable improvement of approximately 1% of AuPRC for a subset of targets and a large improvement of around 4% of AuPRC for the truly OOD target PICdb.
@@ 4.2 The performance gains are largest for the most OOD domains
In figure˜5, for each anchor model and $\gamma>1$, we display the minimal number of patient IDs necessary for this median performance to match the anchor model’s OOD performance. We linearly interpolate $\log(\mathrm{num.patients})\sim\mathrm{performance}$ in between patient numbers. We use this value of required patients to measure dissimilarity: The fewer patients needed from the target domain, the the less value the large external dataset has for that task on the target, and thus the more dissimilar the target domain is. We expect anchor regression to improve performance for the target domains with the largest shifts, effectively lifting the minimal number of patients required to match performance.
@@
We present equivalent figures for the remaining tasks in figures˜19 and 20 in appendix˜C. The results are similar.

### v2

@@ Abstract
The performance of predictive models in clinical settings often degrades when deployed in new hospitals due to distribution shifts. This paper presents a large-scale study of causality-inspired domain generalization on heterogeneous multi-center intensive care unit (ICU) data. We apply anchor regression and introduce anchor boosting, a novel, tree-based nonlinear extension, to a large dataset comprising 400,000 patients from nine distinct ICU databases. We find that anchor regularization yields improvements of out-of-distribution performance, particularly for the most dissimilar target domains. The methods appear robust to violations of theoretical assumptions, such as anchor exogeneity. Furthermore, we propose a novel conceptual framework to quantify the utility of large external data datasets. By evaluating performance as a function of available target-domain data, we identify three regimes: (i) a domain generalization regime, where only the external model should be used, (ii) a domain adaptation regime, where refitting the external model is optimal, and (iii) a data-rich regime, where external data provides no additional value.
@@
Accepted for publication in RSS: Data Science and Artificial Intelligence.
@@ 1 Introduction
A standard assumption in predictive modeling is that training and test data come from the same distribution. This assumption often fails in real-world scenarios. For example, in clinical applications, test data may originate from a different time period or hospital than the training data. When these distribution shifts occur, model performance tends to drop significantly (Barak-Corren et al., 2021; Guo et al., 2021; Roland et al., 2022; Yang et al., 2022; van de Water et al., 2024; Hüser et al., 2024).
@@
In this paper, we apply and extend anchor regression for medical prediction. Recognizing that linear models may be insufficient to capture the complex feature interactions in clinical data, we first propose a novel, nonlinear extension to anchor regression based on gradient boosting trees (Friedman, 2001), popular in clinical predictive modeling (Hyland et al., 2020; Lyu et al., 2024; Hüser et al., 2024). We then conduct a large-scale empirical study, applying linear anchor regression and our nonlinear extension to predict adverse events in intensive care units (ICUs). We assess the method’s ability to improve distributional robustness on a strongly heterogeneous dataset aggregating nine ICU databases (Burger et al., 2024; Burger et al., 2025). Finally, we propose a conceptual framework for quantifying the utility of large external datasets, particularly when target domain data is scarce.
@@ 1.1 Our contribution
We conduct the largest-scale application of anchor regression (Rothenhäusler et al., 2021) and its variants to date, using a dataset of 400,000 patients and 10 million observations from nine distinct ICU databases. To our knowledge, this work is also the first application of anchor regression and the largest application of a causality-inspired method to a medical prediction problem. In a setting where many other domain generalization methods have been shown to provide little benefit over simple baselines (Rockenschaub et al., 2024), we show that the anchor methods yield notable performance improvements, particularly for the most out-of-distribution target domains.
@@ 2 Data description
We use data from Burger et al. (2024); Burger et al. (2025) who build upon the R-package ricu (Bennett et al., 2023) to harmonize and aggregate ICU data from different sources. We will describe this dataset in further detail next.
@@ 2.1 ICU datasets included
We focus on the following 9 ICU datasets: The eICU Collaborative Research Database (Pollard et al., 2018) is a multi-center critical care database of $207$ hospitals throughout the USA. It contains data from 188’257 patient stays from 2015 and 2016. The Medical Information Mart for Intensive Care (MIMIC) III (Johnson et al., 2016) and IV (Johnson et al., 2023) contain data from the Beth Israel Deaconess Medical Center (BIDMC) in Boston (USA). The BIDMC switched critical care information systems in 2008 from Philips CareVue to iMDsoft Metavision. We consider only the CareVue subset of MIMIC-III to avoid an overlap with MIMIC-IV. This MIMIC-III CareVue subset contains data from 34’154 patient stays from 2001–2008 and includes neonatal patient stays. MIMIC-IV contains data from 93’679 patient stays from 2008–2022. The Northwestern ICU database (NWICU) (Moukheiber et al., 2024) is a multi-center critical care database of 12 hospitals around Chicago (USA). It contains data from 28’150 patient stays from 2020–2022. The High-Resolution ICU dataset (HiRID) (Hyland et al., 2020), Amsterdam University Medical Center database (AUMCdb) (Thoral et al., 2021), and Salzburg Intensive Care database (SICdb) (Rodemund et al., 2024), are single-center European datasets. HiRID contains data from 33’586 patient stays from 2008–2016. AUMCdb contains data from 22’897 patient stays from 2003–2016. SICdb contains data from 27’115 patient stays from 2013–2021. The Paediatric Intensive Care database (PICdb) (Li et al., 2019) contains data from the Children’s Hospital of Zhejiang University School of Medicine in China. It contains data from 13’516 patient stays from 2010–2018. The Critical Care Database Comprising Patients With Infection at Zigong Fourth People’s Hospital (Xu et al., 2022) in China contains data of patients with a suspected infection. It contains data from 2’583 patient stays from 2019 and 2020. All datasets except for AUMCdb are available on PhysioNet (Goldberger et al., 2000). We summarize these data in table˜2 in appendix A.
@@ 2.3 Patient outcomes
We visualize the resulting outcome distributions in figure˜1 and provide summary statistics in table˜2 in appendix A. See also github.com/eth-mds/icu-features.
@@ 2.4 Feature engineering and pre-selection of variables
The hourly time-series data, extracted by Burger et al. (2024); Burger et al. (2025) using ricu (Bennett et al., 2023) must be transformed into a static feature set to be used with standard linear or tree-based methods. Following the literature, we engineer features over a backwards looking time window to capture some temporal dynamics of a patient’s history. We use an 8 hour horizon for circulatory failure and log(lactate) prediction and a 24 hour horizon for acute kidney injury and log(creatinine) prediction. These features include a missingness indicator, the mean, standard deviation, maximum, minimum, and a linear trend for continuous variables, the mode and a missingness indicator for categorical variables, and an indicator and the average rate for treatments. See github.com/eth-mds/icu-features and appendix A.2 for details.
For computational reasons and to reduce model complexity, we do not use all available variables of Burger et al. (2024); Burger et al. (2025)’s export in our models. For the prediction of circulatory failure and log(lactate), we use the top 20 variables of Hyland et al. (2020) according to their table 1. For the prediction of acute kidney injury and log(creatinine), we use the top variables of Lyu et al. (2024) according to their figure 8a. See appendix A.3 for details. Together with feature engineering, this results in 100–200 covariates for each task.
@@ 2.5 Sources of heterogeneity
Because of the cohort selection and variable availability, we expect the largest differences in cohort composition and variable availability for NWICU, PICdb, and Zigong. We therefore designate these as “truly out-of-distribution” datasets. In section˜4, except for sections˜4.6 and 4.7, we will use these only for evaluation and not for training. We believe that this reflects a realistic use-case, where a practitioner targeting a general adult ICU would likely exclude data from very dissimilar cohorts, such as the pediatric population in PICdb, from the training set.
@@ 3 Methods
This section details our methodological approach. We begin by introducing linear anchor regression (Rothenhäusler et al., 2021). We then propose anchor boosting, a novel, nonlinear extension based on gradient boosting trees. Finally, we describe our refitting procedure and introduce a taxonomy to quantify the value of external data.
@@ 3.1 Linear anchor regression
Consider a discrete anchor variable given by environments $e\in{\cal E}$, for example, the ICU datasets ${\cal E}=\{\mathrm{AUMCdb},\ \mathrm{eICU},\ \ldots,\ \mathrm{SICdb}\}$. The linear anchor regression estimator is defined as
@@
where $\gamma\geqslant 1$ is the invariance regularization parameter and $n_{e}$ are the number of samples from environment $e$. That is, anchor regression penalizes differences in the environments’ mean residuals. From a theoretical perspective, the anchor regression estimator optimizes a worst-case risk over a set of new, unseen environments $e\notin{\cal E}$, for example $e=\mathrm{PICdb}$, with distribution shifts similar to the heterogeneity seen within the training data, but of a larger magnitude scaled by $\gamma$.
@@ 3.2 Nonlinear anchor boosting (regression)
The theoretical robustness guarantees of anchor regression mentioned above apply for a linear model (Rothenhäusler et al., 2021) or when using a nonlinear embedding (Sola et al., 2025). Such guarantees are unavailable for the proposed anchor boosting, except for some very preliminary result on risk constancy with full invariance corresponding to $\gamma=\infty$ (Bühlmann, 2020, proposition 5.1). Nonlinear methods generally cannot reliably extrapolate to shifted regimes (Christiansen et al., 2022). While linear models may diverge in such settings (correctly, if the model is true), tree ensembles inherently predict constant values outside the training range. This keeps extrapolation of prediction in the range of observed values, thereby handling distribution shifts conservatively. Relying on this stability and the intuition from the linear case that homogenizing residuals across environments promotes robustness, we optimize equations˜1 and 2 by replacing $X_{i}\beta$ with a nonlinear function $f(X_{i})$.
We focus here on boosted tree learners (Friedman, 2001), a popular method for ICU data (Hyland et al., 2020; Lyu et al., 2024). Analogously to (2), let
@@
be the anchor loss with gradient and Hessian with respect to $f$ as follows:
$g(f,y)=-(y-f)-(\gamma-1)P_{A}(y-f)\quad\text{and}\quad H(f,y)=\mathrm{Id}+(\gamma-1)P_{A}.$
Following the construction of gradient boosting (Friedman, 2001), let $\hat{f}^{j}$ be the boosted learner after $j$ steps of boosting. We initialize with $\hat{f}^{0}:=\frac{1}{n}\sum_{i=1}^{n}y_{i}$. We then fit the negative gradient against $X$ using a decision tree $\hat{t}^{j+1}:=-g(\hat{f}^{j}(X),y)\sim X$. Let $M\in{\mathbb{R}}^{n\times\mathrm{num.\ leafs}}$ be the one-hot encoding of $\hat{t}^{j+1}(X)$’s leaf node indices. Then, $M^{T}\,g(\hat{f}^{j}(X),y)$ and $M^{T}\,H(\hat{f}^{j}(X),y)\,M$ are the gradient and Hessian of the loss function $\ell(\hat{f}^{j}(X)+\hat{t}^{j+1}(X),y)=\ell(\hat{f}^{j}(X)+M\hat{\beta}^{j+1},y)$ with respect to $\hat{t}^{j+1}$’s leaf node values $\hat{\beta}^{j+1}\in{\mathbb{R}}^{\mathrm{num.\ leafs}}$. We set them using a second order optimization step to
$\hat{\beta}^{j+1}=-\left(M^{T}\,H(\hat{f}^{j}(X),y)\,M\right)^{-1}M^{T}\,g(\hat{f}^{j}(X),y).$
@@
A nonlinear extension of anchor regression based on boosting was already proposed by Bühlmann (2020). However, this proposal does not use second order optimization to update the leaf node values, which appears to be crucial when using larger values of $\gamma$, especially for classification as described in section˜3.3. See appendix B.3 for details. In related work, Ulmer et al. (2025) fit a random forest to a linear rescaling of the data. In Ulmer and Scheidegger (2025) they discuss how this can be applied to anchor regression, resulting in split points for each tree chosen to directly minimize the anchor loss. However, this approach increases computational complexity beyond a near-linear relationship with sample size, making it prohibitive for the large datasets that we address in this paper.
@@ 3.3 Nonlinear anchor boosting (classification)
We apply anchor regularization to binary classification tasks. Kook et al. (2022) suggest to use the gradient of the log-likelihood as score residuals. We use a probit link function, as, in contrast to the logistic link, the resulting anchor classification objective is convex. This convexity is necessary for stable second-order updates of the tree-leaf values. See appendix B.2 for details.
Write $\Phi$ and $\varphi$ for the Gaussian distribution’s cumulative distribution function and probability density function. For binary classification with scores $f\in{\mathbb{R}}^{n}$ and $y\in\{-1,1\}^{n}$, the negative log-likelihood is $\sum_{i=1}^{n}\log(\Phi(y_{i}f_{i}))$ with negative gradient $r:=y\cdot\varphi(f)/\Phi(yf)$. We thus apply the same procedure as in section˜3.2 to the loss
@@
replacing the squared error $\frac{1}{2}(y_{i}-f_{i})^{2}$ in equation˜3 with the negative log-likelihood $\log(\Phi(y_{i}f_{i}))$ and the residuals $y_{i}-f_{i}$ with the likelihood score residuals $r_{i}$. We initialize with $\hat{f}^{0}=\Phi^{-1}(\frac{1}{n}\sum_{i=1}^{n}\frac{y_{i}+1}{2})$. See appendix B.1 for derivations of the gradient and Hessian of equation˜4.
@@ 4.1 Anchor regularization improves generalization to some ICU datasets
Figure 3: Anchor boosting’s OOD MSE predicting log(creatinine) as a function of $\gamma$, using one-hot-encoded dataset ID as anchor. See section˜4.3 for details on the LOEO-CV model selection.
@@
Figure˜17 in appendix C shows anchor boosting’s OOD area under the precision-recall curve (AuPRC) predicting acute kidney injury within 48 hours, the classification task corresponding to log(creatinine) regression. AuPRC is a common metric for predicting rare events on ICU data. Anchor regularization with $\gamma>1$ yields a considerable improvement of approximately 1% of AuPRC for a subset of targets and a large improvement of around 4% of AuPRC for the truly OOD target PICdb.
@@ 4.2 The performance gains are largest for the most OOD domains
In figure˜5, for each anchor model and $\gamma>1$, we display the minimal number of patient IDs necessary for this median performance to match the anchor model’s OOD performance. We linearly interpolate $\log(\mathrm{num.patients})\sim\mathrm{performance}$ in between patient numbers. We use this value of required patients to measure dissimilarity: The fewer patients needed from the target domain, the less value the large external dataset has for that task on the target, and thus the more dissimilar the target domain is. We expect anchor regression to improve performance for the target domains with the largest shifts, effectively lifting the minimal number of patients required to match performance.
@@
We present equivalent figures for the remaining tasks in figures˜19 and 20 in appendix C. The results are similar.